OpenStudy (jessiegonzales):

MEDAL!!!! 2. What is the order of the numbers from least to greatest? A= 6.2 * 10^-5 B = 12.3 * 10^-7 C=5.2 * 10^3 D=3.7 * 10^-5

OpenStudy (solomonzelman):

simplify and get an exact value of each option.

OpenStudy (jessiegonzales):

@SolomonZelman IS IT D>A>C>B

OpenStudy (jessiegonzales):

@Hero

OpenStudy (jessiegonzales):

sorry i mean D>C>A>B

OpenStudy (solomonzelman):

I would have to check. Hold on. \(\large\color{midnightblue}{ \rm 6.2 \times 10^{-5}~~=0.000062 }\) \(\large\color{midnightblue}{ \rm 12.3 \times 10^{-7}=0.00000123 }\) \(\large\color{midnightblue}{ \rm 5.2 \times 10^{3}~~~~=5,200 }\) \(\large\color{midnightblue}{ \rm 3.7 \times 10^{-5}~~=0.000037 }\)

OpenStudy (solomonzelman):

Can you now write it in correct order ?

OpenStudy (solomonzelman):

In B it might be 1/10th greater, but it's either way the smallest value.